OpenStudy (anonymous):

Find the hexadecimal expansion of (ABC)16 + (2F5)16 (Ans: (DB1)16

OpenStudy (mathmate):

Add digit by digit: ABC + DB1 ------ 186D C+1=D B+B=11+11=22=6 with carry A+D=10+13+carry=24(base 10)=18 (hex)

OpenStudy (anonymous):

how is (b+b)=22?

OpenStudy (mathmate):

Sorry, let's start over: Add digit by digit: ABC + 2F5 ------ DB1 C+5=12+5=17(base 10) = 1 with carry B+F+carry=11+15+carry=27(base 10)=B with carry A+2+carry=D

OpenStudy (mathmate):

B(hex)=11(10) So 11(10)+11(10)=22(10)=16(hex)

OpenStudy (anonymous):

You are the king !!

OpenStudy (mathmate):

Thank you! I hope you can tackle anything now!